# Anomaly Induced Inflation in the Minimal Quartic Extension of Einstein’s Gravity

### v1

@@ Abstract
A particular extension of Einstein’s General Relativity up to and including quartic terms in the curvature tensor is minimal in the sense that it has a unique maximally symmetric vacuum and only a massless spin-2 excitation in its spectrum around the vacuum. We study the inflation phase of the universe in the minimal quartic extension of Einstein’s gravity in the presence of trace anomaly terms in the Standard Model and the Minimal Supersymmetric Standard Model and show that the theory allows a quasi-de Sitter phase with a sufficient number of $e$-foldings and is compatible with the spectral index and the tensor to scalar ratio.
@@ 1 Introduction
The extreme homogeneity and the isotropy in the Cosmic Microwave Background (CMB) data [1] can be understood if the universe had gone through a brief period of accelerated expansion phase, or if the universe were a lot older. The latter option is unlikely. Within General Relativity such an expansion requires the existence of additional scalar fields, such as ”the inflaton” that first inflates the universe, gets affected by the inflation, and creates friction to gracefully exit the inflation phase. Another possible path to inflation is to modify the underlying gravity theory as was done by Starobinsky [2]. Here we follow this second route and study the inflationary phase of a modified gravity theory built on the curvature tensors up to and including the quartic terms. The theory was initially obtained as a subset of Born-Infeld-type gravity theory that was judiciously constructed to have all the good features of Einstein’s theory, such as the existence of a unique maximally symmetric vacuum and the non-existence of degrees of freedom besides a massless spin-2 graviton, but is much better behaved in the ultraviolet region of gravity. Here we first describe the theory before we study its cosmological solutions.
@@
The Born-Infeld (BI) type gravity theories are a class of theoretical frameworks in gravitational physics that extend Einstein’s General Theory of Relativity (GR). They emerge from the need to address the inherent singularities present in GR, such as those associated with black holes and the Big Bang, by introducing higher derivative, and higher curvature corrections to the gravitational action. These corrections are inspired by the Born-Infeld electrodynamics (to be historically accurate these theories were first considered by Schrodinger and later by Eddington in gravity), originally formulated to regularize the self-energy of point charges. BI gravity theories have been studied in the context of string theory and modified gravity theories, offering new perspectives on the fundamental nature of gravity and its behavior in extreme conditions. The general form of the action of the BI type gravity theories can be given by
@@
where $\gamma$ is the Born-Infeld parameter and $A_{\mu\nu}$ is a second rank tensor constructed from the curvatures. Of course a proper $A_{\mu\nu}$ must be found that is consistent with all known tests of gravity and that also does not lead to inconsistencies, such as ghosts in the theory. For example, the naive, obvious choice $A_{\mu\nu}=R_{\mu\nu}$ leads to massive spin-2 ghosts in the spectrum. How ghosts can be avoided in these types of theories is a rather complicated problem, especially about a de-Sitter or anti-de-Sitter background. But fortunately, this problem has been fully solved which we discuss below.
@@
$I=\frac{1}{2\kappa_{0}\gamma}\int d^{4}x\left(\sqrt{-\operatorname{det}\left(g_{\mu\nu}+4\gamma A_{\mu\nu}\right)}-\left(4\gamma\Lambda_{0}+1\right)\sqrt{-\operatorname{det}g}\right)$ (2)
@@
Here, $R_{\mu\nu}$ is the Ricci tensor while $S_{\mu\nu}$ is the traceless Ricci tensor, and $C_{\mu\rho\nu\sigma}$ is the Weyl tensor. The features, formulation of this theory, and its extension to generic $n$ dimensions were laid out in detail in [4], which relies on the foundation given in [5]. Note that this theory is in a sense the minimal BI gravity extension of Einstein’s theory as one can add more powers of curvature inside the determinant to obtain non-minimal extensions. Note also that as far as these theories are concerned, the $2+1$ dimensional BI gravity is rather unique, describing a massive graviton (instead of a massless one) with a remarkably simple Lagrangian of the form ${\mathcal{L}}=\sqrt{-\det\left(g_{\mu\nu}+\gamma G_{\mu\nu}\right)}$ [6] where $G_{\mu\nu}$ is the Einstein tensor. The success of this particular theory led us to search for similar theories in four and higher dimensions. Let us recapitulate some important aspects of the theory (2) that arise from the properties of the determinant and are not shared by most higher derivative gravity theories:
@@
  • When expanded in small curvature ($|\gamma R|\ll 1$), the theory reproduces General Relativity as demanded by low energy observations at the first order; and at the quadratic order, it reproduces GR modified with the Gauss-Bonnet term, which at the classical level is just GR due to the topological nature of the Gauss-Bonnet term in four dimensions. The specific cubic and quartic order theories that we shall work with here were studied recently in [8] in the context of black hole physics.
2 Cosmological solutions
@@
It is clear that for the RW metric, the Weyl tensor vanishes identically and hence the parameter $a$ does not contribute to the reduced action. Note also that for the particular case $c=-1$ and $b=-5/2$, one has a purely quartic theory since $X_{1}=X_{2}$ of which the spherically symmetric black hole solutions were studied in [8]. Equipped with this, the necessary and sufficient equation to solve is the one coming from
@@
irrespective of the presence of a bare cosmological constant. The de Sitter solution can be found using the first Friedmann equation in (14).
@@
where we defined an effective cosmological constant $\Lambda=3H_{0}^{2}$ and corresponding dimensionless parameters $\lambda=\gamma\Lambda$ and $\lambda_{0}=\gamma\Lambda_{0}$ for effective and bare one respectively. Note that the last equation is (4). This polynomial equation has four solutions, two of which are imaginary and one of them is not physical in the sense that it gives $\lambda>1/4$ which corresponds to $\kappa<0$ according to (5) so that we have a unique de Sitter solution for given $\lambda_{0}$.
Here, we provide some remarks concerning the free parameters and the higher-order curvature terms within the Lagrangian. As mentioned previously, setting $a=0$, $b=-5/2$, and $c=-1$, the theory assumes a purely quartic nature, resulting in a Lagrangian with a finite number of curvature terms, specifically up to and including $R^{4}$ order terms. Alternatively, if we opt to keep these parameters as free variables, the conventional approach to derive field equations involves expanding $\operatorname{det}(\delta^{\nu}_{\mu}+4\gamma A_{\mu}^{\nu})$ and its square root in the small $\gamma$ limit. Consequently, the Lagrangian incorporates an infinite number of higher-order terms. However, in our analysis we found that the parameters $a$, $b$ and $c$ do not appear in the de Sitter equation and terms beyond the fourth order in curvature exhibit no influence on both the de Sitter solution and the linear perturbations around it. This observation implies that, when considering the simplified cosmological scenarios of a de Sitter universe and the quasi-de Sitter analysis, it is sufficient to focus on the lower-order curvature terms, thereby we can safely set the parameters to the desired values. In that case, the Lagrangian in (2) can be reduced to the simpler form.
@@
This is the minimal quartic extension of Einstein’s gravity in the sense that it has the same particle spectrum and a unique vacuum. With our parameter settings, the explicit form of Lagrangian (18) shows that it is quartic in curvature. Henceforth, in the remainder of the analysis, we shall adopt this simpler theory. In principle, one would like to study the process of inflation within the BI gravity described so far, but it turns out the theory does not have a quasi-de sitter phase. Such an attempt was made by one of us in [12] expecting that the theory would start from the unstable $\kappa<0$, vacuum, that is the repulsive gravity phase, and end up in the $\kappa>0$ vacuum in an inflated state. But, due to a computational error in that work, the claimed result is not correct. Hence, we must resort to the trace anomaly to realize inflation. We describe some basics of the trace anomaly in the next section before applying it to quartic theory.
@@ 3.1 Trace anomaly
where $\beta$s are renormalization group $\beta$-functions that depend on the number of fields present in the theory and given as [14, 15]
@@
$\left\langle T_{\mu}^{\mu}\right\rangle=-\frac{2}{\sqrt{-g}}g_{\mu\nu}\frac{\delta\bar{\Gamma}^{(1)}}{\delta g_{\mu\nu}}=\alpha W-\beta\mathcal{G}+\delta\square R\,,$ (28)
where $\alpha=\beta_{1}$, $\beta=\beta_{2}$, and $\delta=\beta_{3}$ for global conformal symmetry. Here, the emergence of the $\Box R$-term in the trace of the energy-momentum tensor is especially important since this is the term responsible for the Starobinsky instability. We should note that the coefficients $\beta_{1}$ and $\beta_{2}$ are independent of the renormalization scheme, but $\beta_{3}$ is not. For example, the notion adopted here that is the result given by the zeta-function regularization or predicted by AdS/CFT gives $-18$ as the coefficient of $N_{V}$ in the expression of $\beta_{3}$ while the result given by dimensional regularization has $+12$. The renormalization-dependent nature of $\delta$ suggests the possibility of adding a finite $R^{2}$ counterterm to the action at the classical level. When $\mathcal{N}=4$ super Yang-Mills theory is considered, for example, the field content can be given by a single parameter $N$ which is a large number. In that case, the effect of $\Box R$ vanishes and then the inflation never ends. However, $\delta$ can be adjusted to any desired value by adding the finite counterterm.
@@
This counter-term explicitly breaks conformal invariance. In the literature, there is an ambiguity in the parameter $\delta$ and a comprehensive examination of this matter is available in [14, 16, 17]. The presence of $S_{ct}$ implies that we are effectively dealing with a higher derivative theory of gravity but it is arbitrary to consider it as part of the gravitational action or as part of the matter action. Since we are using the Born-Infeld action as the gravitational sector, the origin of the parameter $\delta$ can be regarded as a matter contribution. Hence, we provide particle content of three example models in Table 1, and adopted the Minimum Standard Model in our analysis mainly.
@@
Finally, the scenario can be succinctly encapsulated in the following manner: In curved space-time, the action governing the massless matter fields, that are scalars, Dirac spinors, and vectors, exhibits conformal invariance, yet the presence of one-loop vacuum contributions introduces certain divergences. Within the context of renormalization, specific counterterms are introduced to remove the singularities within the divergent part, albeit at the cost of breaking the conformal invariance inherent to the matter action itself. Classically, within a conformally invariant theory, the trace of the energy-momentum tensor is expected to be zero. However, the renormalization process results in the emergence of an anomalous trace of the energy-momentum tensor in (28), commonly referred to as the conformal anomaly or the quantum anomaly.
3.2 Trace anomaly and minimal quartic theory
@@
where $\mathcal{F}$ is defined in (20), and $\mathcal{F}_{R}$ and $\mathcal{F}_{\mathcal{G}}$ are partial derivative of $\mathcal{F}(R,\mathcal{G})$ with respect to $R$ and $\mathcal{G}$. The energy-momentum tensor $\left\langle T_{\mu\nu}\right\rangle=\left\langle 0\left|T_{\mu\nu}\right|0\right\rangle$ emerges from the vacuum expectation value of the quantum corrections. Then, by taking trace of the field equations, we have
@@
with $W=0$ in flat RW metric. Using the conservation law in Eq. (35) and eliminating the pressure, we can write
$\frac{d}{dt}\left(\rho a^{4}\right)=-\dot{a}a^{3}\left(-\rho+3p\right)=\equiv a^{3}\dot{a}\left(24\beta\frac{\dot{a}^{2}\ddot{a}}{a^{3}}+\delta(\ddot{R}+3H\dot{R})\right).$ (37)
@@
$0\displaystyle 324H_{0}^{8}\gamma^{3}+108H_{0}^{6}\gamma^{2}+6H_{0}^{4}\kappa_{0}\beta-3H_{0}^{2}+\Lambda_{0}=$ 0 (41)
$\frac{2\beta_{0}\lambda^{2}}{3}+4\lambda^{4}+4\lambda^{3}-\lambda+\lambda_{0}=0$
@@
3.4 Dynamics of inflation
We first analyze our main differential equation (40) with a phase portrait analysis. Using a change of variables,
$x=H(t)\quad\text{and}\quad y=\dot{H}(t),$ (42)
it can be written as a first order differential equation
With this reparametrization, the phase portrait of the system can be analyzed in the $xy$-plane as in Fig. 1. We provided the phase portrait for the Minimum Standard Model (MSM) in four different regimes.
In the first scenario depicted in Fig. 1 (a), we examine the case where there is no contribution from the modified gravity, representing a pure trace-anomaly situation within the framework of Einstein’s gravity. Two critical points can be seen along the $y=0$ axis, corresponding to two distinct de Sitter solutions. For the particle content of MSM, the second solution is identified as unstable. Consequently, the universe is presumed to have originated nearly from this point, having a negative $\dot{H}$, subsequently transitioning towards the first critical point, which acts as the final attractor of the system, determined by $\Lambda_{0}$, and may correspond to the late-time acceleration. In the absence of an intrinsic cosmological constant, the ultimate attractor point resides at $(x,y)=(0,0)$. Thus, during the inflationary epoch, the Hubble parameter $H$ experiences a slow decrease while maintaining $\dot{H}$ in near constancy, eventually oscillating around the final attractor, indicative of the reheating era.
(a) Einstein’s gravity with trace-anomaly. ($\gamma=0$)
(b) Minimal quartic model without trace-anomaly. ($\gamma=0.0025$)
(c) Minimal quartic model with trace-anomaly. ($\gamma=0.0025$)
(d) Minimal quartic model with trace-anomaly. ($\gamma=0.0060$)
Figure 1: The phase-portrait of the system described by the EoM 40 for the MSM case with parameter settings $x=H(t)$ and $y=\dot{H}(t)$. The figure is plotted in $c=\hbar=\kappa_{0}=1$ units and the value of the bare cosmological constant is chosen arbitrarily as $\Lambda_{0}=6$.
In the second scenario, depicted in Fig. 1 (b), we switch off the trace anomaly and investigate the dynamics of pure modified gravity. Once again, two distinct de Sitter solutions are evident along the $y=0$ axis. Additionally, a solid black line is introduced along the x-axis to indicate a non-physical region where $\lambda>1/4$, a characteristic feature typical of Born-Infeld-type theories, setting an upper limit on the effective cosmological constant. Consequently, the second solution is nonphysical, as it intersects with the solid line, resulting in the existence of a unique de Sitter solution. In this context, a critical surface manifests along the $y=0$ axis, changing the characteristics of critical points both below and above this surface. The unique de Sitter solution remains stable within the domain where $\dot{H}<0$, yet transitions to instability when $\dot{H}>0$ is encountered. Furthermore, the critical surface is identified as unstable within these points, while exhibiting attractor behavior beyond this region.
In Fig. 1 (c), we explore our modified gravity theory gently introducing a small value for $\gamma$ in the presence of the trace anomaly. In comparison to the first scenario, the introduction of modified gravity results in the emergence of three critical points situated above the de Sitter solutions observed in the original trace-anomaly case. However, one of these critical points exhibits a notably high $\dot{H}$ value, surpassing the upper limit of this plot and consequently playing no role in the inflationary solutions. The remaining two critical points reside above the critical points along the $y=0$ axis and are positioned on a critical surface represented by a curve that converges toward the x-axis as $H\rightarrow\infty$. Notably, the critical surface with critical points on it in pure modified gravity case is raised within the $\dot{H}>0$ region, while still maintaining the de Sitter solution due to the trace anomaly within the physically plausible range, $\kappa>0$. Under an initial condition where $\dot{H}\lesssim 0$, the dynamics of inflation remain largely unaffected by the presence of modified gravity.
In Fig. 1 (d), to understand the effect of modified gravity on trace-anomaly-driven inflation, we play with the value of $\gamma$ and increase the intensity it. Increasing it makes the above critical points closer to the de Sitter points, while the solid line indicating $\lambda>0$ region, started to cover up the critical point corresponding to inflation so that inflation cannot happen. This imposes an upper bound on $\gamma$ so that the effects of modified gravity should not exceed the trace anomaly for a viable inflationary model.
Furthermore, we investigate the scenario wherein the sign of the $\Box R$-term is positive, indicative of an initial stable point. This condition finds representation in the context of MSSM as an illustrative example model encompassing the particle content pertinent to this case. In Fig.2, the phase-portrait of the system is plotted without and with the effects of modified gravity for the MSSM case respectively, and from Fig.(a), it can be seen that the second de Sitter solution due to the trace-anomaly is a stable critical point. In this case, a graceful exit from Inflation requires its mechanism to disrupt this state. In Ref. [18], a possible transition from stable to unstable phases of inflation is discussed, and different quantum mechanisms that may produce large $R^{2}$-term during the inflation are proposed for a graceful exit for example. On the contrary, we show that the minimal extension of Einstein’s gravity makes possible exit from inflation without any other mechanism. In Fig.2(b), we have plotted the contribution of modified gravity in this case. As in MSM case, modified gravity creates additional critical points on the phase-portrait, visible ones of those in the plotted region are labeled as $P_{i}(x_{i},y_{i})$ in the graph, but this time the two of them lying on a critical surface are in $\dot{H}<0$ region. Note that increasing the intensity of modified gravity makes this critical surface closer to the $y=0$ line. Therefore, if the universe starts with $H\simeq x_{1}\simeq x_{2}$ and a small value of $\dot{H}$ from below this surface, $\dot{H}\lesssim y_{2}$ for example, then the exit from inflation can occur due to the attractor behavior of the surface between 00 and $x_{3}$. This prevents $\dot{H}$ from flowing through negative infinity and ends inflation with $H\lesssim x_{3}$. If the strength of modified gravity is high enough, the initial point for $\dot{H}$ can be in the slow-roll regime, i.e. $\left|\dot{H}/H^{2}\right|\ll 1$ while keeping $\kappa$ positive.
(a) Trace-anomaly in Einstein’s gravity with MSSM particle content
Figure 2: The phase-portrait of the system described by the EoM 40 for MSSM case. The figure is plotted in $c=\hbar=\kappa_{0}=1$ units and the parameters are set as $\gamma=0.004$ and $\Lambda_{0}=6$.

### v2

@@ Abstract
A particular extension of Einstein’s General Relativity up to and including quartic terms in the curvature tensor is minimal in the sense that it has a unique maximally symmetric vacuum and only a massless spin-2 excitation in its spectrum around the vacuum. We study the inflation phase of the universe in this minimal quartic extension of Einstein’s gravity in the presence of trace anomaly terms coming from the Standard Model fields and the fields of the Minimal Supersymmetric Standard Model. We show that the theory allows a quasi-de Sitter phase with sufficient $e$-foldings.
@@ 1 Introduction
The extreme homogeneity and the isotropy in the Cosmic Microwave Background (CMB) data [1] can be understood if the universe had gone through a brief period of accelerated expansion phase, or if the universe were a lot older. The latter option is unlikely. Within General Relativity such an expansion requires the existence of additional scalar fields, such as the ”inflaton” that first inflates the universe, gets affected by the inflation, and creates friction to exit the inflation phase gracefully. Another possible path to inflation is to modify the underlying gravity theory. Here we follow this second route and study the inflationary phase of a modified gravity theory built on the powers of the curvature tensors up to and including the quartic terms. The theory that we shall study here was initially obtained as a subset of Born-Infeld-type gravity theory that was judiciously constructed to have all the good features of Einstein’s theory, such as the existence of a unique maximally symmetric vacuum and the non-existence of degrees of freedom besides a massless spin-2 graviton, but is much better behaved in the ultraviolet region of gravity. With properly chosen free parameters, Born-Infeld gravity was shown to reduce to a quartic theory in the form of $\mathcal{F}(R,\mathcal{G})$ gravity theories where $\mathcal{G}$ stands for Gauss-Bonnet scalar. This particular quartic theory has a unique de Sitter vacuum, but it does not have an inflationary phase, that is it does not have a quasi-de Sitter solution with enough $e$-foldings. To remedy this, we shall resort to the conformal anomaly that produces additional degrees of freedom besides the massless spin-2 graviton. This scenario was originally proposed by Starobinsky in 1980 [2]. We shall adopt it to our theory, in which, inflation is driven by a quantum anomaly arising from the matter fields’ renormalization procedure. Renormalization results in an anomalous energy-momentum tensor that could generate an unstable vacuum with a higher energy density so that the Universe makes a transition from this into the true vacuum state by inflating from a microscopic size to a macroscopic size.
The paper is organized as follows. In the next section, we first describe the full Born-Infeld gravity theory and its cosmological solutions; we show how the theory reduces to a quartic extension (that is the action is a specific polynomial of the curvature tensor up to the fourth power of the Riemann curvature tensor and its contractions) of Einstein’s theory. In Section 3, we introduce and summarize the idea of trace anomaly before embedding it into our theory. Then, we derive the field equations and de Sitter solutions in the presence of trace anomaly. The analysis continues with the linear perturbation around the de Sitter solutions to find the stability conditions. To study the inflationary dynamics, we first examine the phase portrait of the dynamical system to see its general behavior, and then solve the system numerically for particular inflationary solutions. Lastly, by proposing an approximate solution, we study the matter perturbations during inflation.
2 The minimal extension of Einstein’s gravity
@@
The Born-Infeld (BI) type gravity theories are a class of theoretical frameworks in gravitational physics that extend Einstein’s General Theory of Relativity (GR). They emerge from the need to address the inherent singularities present in GR, such as those associated with black holes and the Big Bang, by introducing higher-derivative and higher-curvature corrections to the gravitational action. These corrections are inspired by the Born-Infeld electrodynamics (to be historically accurate these theories were first considered by Schrodinger and later by Eddington in gravity), originally formulated to regularize the self-energy of point charges. BI gravity theories have been studied in the context of string theory and modified gravity theories, offering new perspectives on the fundamental nature of gravity and its behavior in extreme conditions. The general form of the action of the BI-type gravity theories can be given by
@@
where $\gamma$ is the Born-Infeld parameter and $A_{\mu\nu}$ is a second rank tensor constructed from the curvatures. Of course, a proper $A_{\mu\nu}$ must be found that is consistent with all the known tests of gravity and, that, also does not lead to inconsistencies, such as ghosts in the theory. For example, the naive, obvious choice $A_{\mu\nu}=R_{\mu\nu}$ leads to massive spin-2 ghosts in the spectrum. How ghosts can be avoided in these types of theories is a rather complicated problem, especially about a de-Sitter or anti-de-Sitter background. But fortunately, this problem has been fully solved which we discuss below.
@@
$I=\frac{1}{2\kappa_{0}\gamma}\int d^{4}x\left(\sqrt{-\operatorname{det}\left(g_{\mu\nu}+4\gamma A_{\mu\nu}\right)}-\left(4\gamma\Lambda_{0}+1\right)\sqrt{-g}\right),$ (2)
@@
Here, $R_{\mu\nu}$ is the Ricci tensor while $S_{\mu\nu}$ is the traceless Ricci tensor, and $C_{\mu\rho\nu\sigma}$ is the Weyl tensor. The features, formulation of this theory, and its extension to generic $n$ dimensions were laid out in detail in [4], which relies on the foundation given in [5]. Note that this theory is in some sense the minimal BI gravity extension of Einstein’s theory as one can add more powers of curvature inside the determinant to obtain non-minimal extensions. Note also that as far as these theories are concerned, the $2+1$ dimensional BI gravity is unique, describing a massive graviton (instead of a massless one) with a remarkably simple Lagrangian of the form ${\mathcal{L}}=\sqrt{-\det\left(g_{\mu\nu}+\gamma G_{\mu\nu}\right)}$ [8] where $G_{\mu\nu}$ is the Einstein tensor. The success of this particular theory led us to search for similar theories in four and higher dimensions. Let us recapitulate some important aspects of the theory (2) that arise from the properties of the determinant and are not shared by most higher derivative gravity theories:
@@
  • When expanded in small curvature ($|\gamma R|\ll 1$), the theory reproduces General Relativity as demanded by low energy observations at the first order; and at the quadratic order, it reproduces GR modified with the Gauss-Bonnet term, which at the classical level is just GR due to the topological nature of the Gauss-Bonnet term in four dimensions. The specific cubic and quartic order theories we shall work with here were studied recently in [7] in the context of black hole physics.
2.1 Cosmological solutions and the minimal ‘quartic’ theory
@@
It is clear that for the RW metric, the Weyl tensor vanishes identically; hence the parameter $a$ does not contribute to the reduced action. Note also that for the particular case $c=-1$ and $b=-5/2$, one has a purely quartic theory since $X_{1}=X_{2}$ of which the spherically symmetric black hole solutions were studied in [7]. Equipped with this, the necessary and sufficient equation to solve is the one coming from
@@
irrespective of the presence of a bare cosmological constant. The de Sitter solution can be found using the first Friedmann equation in (14) that yields
@@
where we defined an effective cosmological constant $\Lambda=3H_{0}^{2}$ and corresponding dimensionless parameters $\lambda=\gamma\Lambda$ and $\lambda_{0}=\gamma\Lambda_{0}$ for effective and bare one respectively. Note that the last equation is (4). This polynomial equation has four solutions, two of which are imaginary and one of them is not physical in the sense that it gives $\lambda>1/4$ which corresponds to $\kappa<0$ according to (5) so that we have a unique de Sitter solution for given $\lambda_{0}$. Here, we provide some remarks concerning the free parameters and the higher-order curvature terms within the Lagrangian. As mentioned previously, setting $a=0$, $b=-5/2$, and $c=-1$, the theory assumes a purely quartic nature, resulting in a Lagrangian with a finite number of curvature terms, specifically up to and including $R^{4}$ order terms. Alternatively, if we opt to keep these parameters as free variables, the conventional approach to derive field equations involves expanding $\operatorname{det}(\delta^{\nu}_{\mu}+4\gamma A_{\mu}^{\nu})$ and its square root in the small $\gamma$ limit. Consequently, the Lagrangian incorporates an infinite number of higher-order terms. However, in our analysis, we found that the parameters $a$, $b$ and $c$ do not appear in the de Sitter solution; and terms beyond the fourth order in curvature exhibit no influence on both the de Sitter solution and the linear perturbations around it. This observation implies that, when considering the simplified cosmological scenarios of a de Sitter universe and the quasi-de Sitter analysis, it is sufficient to focus on the lower-order curvature terms, thereby we can safely set the parameters to the desired values. In that case, the Lagrangian in (2) can be reduced to the simpler form  [^1] Note that this is the reduced form of the action (2) with parameters $a$, $b$, $c$ fixed, and this action is not dictated by the trace anomaly that will be introduced in the next section..
@@
This is the minimal quartic extension of Einstein’s gravity in the sense that it has the same particle spectrum and a unique vacuum. With our parameter settings, the explicit form of Lagrangian (18) shows that it is quartic in curvature. Henceforth, in the remainder of the analysis, we shall adopt this simpler theory. In principle, one would like to study the process of inflation within the BI gravity described so far, but it turns out the theory does not have a quasi-de sitter phase. Such an attempt was made by one of us in [12] expecting that the theory would start from the unstable $\kappa<0$, vacuum, that is the repulsive gravity phase, and end up in the $\kappa>0$ vacuum in an inflated state. But, due to a computational error in that work, the claimed result is not correct. Although both Einstein’s gravity and our extended version admit maximally symmetric vacuum solutions, they fail to generate an early inflationary phase on their own. It turns out that the contributions from the matter sector, specifically the trace anomaly arising from the quantum corrections of matter fields in our case, are crucial for realizing an inflationary phase. Without the matter part, our extended theory based solely on the Born-Infeld gravitational action (2) does not provide a viable quasi-de Sitter inflationary solution, much like the ordinary Einstein’s theory without additional fields, $G_{\mu\nu}=0$. Hence, we must resort to the trace anomaly to realize inflation. We describe some basics of the trace anomaly in the next section before applying it to quartic theory.
@@ 3.1 Trace anomaly
where $\beta$s are renormalization group $\beta$-functions that depend on the number of fields present in the theory; and given as [14, 15]
@@
$\left\langle T_{\mu}^{\mu}\right\rangle=-\frac{2}{\sqrt{-g}}g_{\mu\nu}\frac{\delta\bar{\Gamma}^{(1)}}{\delta g_{\mu\nu}}=\alpha^{\prime}W-\beta\mathcal{G}+\delta\square R\,,$ (28)
where $\alpha^{\prime}=\beta_{1}$, $\beta=\beta_{2}$, and $\delta=\beta_{3}$ for global conformal symmetry. Here, the emergence of the $\Box R$-term in the trace of the energy-momentum tensor is important since this is the term responsible for the Starobinsky instability. We should note that the coefficients $\beta_{1}$ and $\beta_{2}$ are independent of the renormalization scheme, but $\beta_{3}$ is not. For example, the notion adopted here that is the result given by the zeta-function regularization or predicted by AdS/CFT gives $-18$ as the coefficient of $N_{V}$ in the expression of $\beta_{3}$ while the result given by dimensional regularization has $+12$. The renormalization-dependent nature of $\delta$ suggests the possibility of adding a finite $R^{2}$ counter-term to the action at the classical level. When $\mathcal{N}=4$ super Yang-Mills theory is considered, for example, the field content can be given by a single parameter $N$ which is a large number. In that case, the effect of $\Box R$ vanishes and then the inflation never ends. However, $\delta$ can be adjusted to any desired value by adding the finite counter-term.
@@
This counter-term explicitly breaks conformal invariance. In the literature, there is an ambiguity in the parameter $\delta$ and a comprehensive examination of this matter is available in [14, 16, 17]. The presence of $S_{ct}$ implies that we are effectively dealing with a higher derivative theory of gravity, but it is arbitrary to consider it as part of the gravitational action or as part of the matter action. Since we are using the Born-Infeld action as the gravitational sector, the origin of the parameter $\delta$ can be regarded as a matter contribution. Hence, we provide particle content of three example models in Table 1, and adopted the Minimum Standard Model in our analysis mainly.
@@
Finally, the scenario can be succinctly encapsulated in the following manner: In curved space-time, the action governing the massless matter fields, that are scalars, Dirac spinors, and vectors, exhibits conformal invariance, yet the presence of one-loop vacuum contributions introduces certain divergences. Within the context of renormalization, specific counter-terms are introduced to remove the singularities within the divergent part, albeit at the cost of breaking the conformal invariance inherent to the matter action itself. Classically, within a conformally invariant theory, the trace of the energy-momentum tensor is expected to be zero. However, the renormalization process results in the emergence of an anomalous trace of the energy-momentum tensor in (28), commonly referred to as the conformal anomaly or the quantum anomaly.
3.2 Trace anomaly within minimal quartic theory
@@
where $\mathcal{F}$ is defined in (20), and $\mathcal{F}_{R}$ and $\mathcal{F}_{\mathcal{G}}$ are partial derivative of $\mathcal{F}(R,\mathcal{G})$ with respect to $R$ and $\mathcal{G}$. The energy-momentum tensor $\left\langle T_{\mu\nu}\right\rangle=\left\langle 0\left|T_{\mu\nu}\right|0\right\rangle$ emerges from the vacuum expectation value of the quantum corrections. For a further discussion on the trace anomaly, see [18] and the references therein.
Then, by taking trace of the field equations, we have
@@
with $W=0$ in flat RW metric. Using the conservation law in (35) and eliminating the pressure, we can write
$\frac{d}{dt}\left(\rho a^{4}\right)=-\dot{a}a^{3}\left(-\rho+3p\right)=a^{3}\dot{a}\left(24\beta\frac{\dot{a}^{2}\ddot{a}}{a^{3}}+\delta(\ddot{R}+3H\dot{R})\right).$ (37)
@@
$324H_{0}^{8}\gamma^{3}+108H_{0}^{6}\gamma^{2}+6H_{0}^{4}\kappa_{0}\beta-3H_{0}^{2}+\Lambda_{0}=0\,,$ (41)
$\frac{2\beta_{0}\lambda^{2}}{3}+4\lambda^{4}+4\lambda^{3}-\lambda+\lambda_{0}=0\,,$